OpenStudy (anonymous):

What type of line is the graph of 4x + y = -3?

563blackghost (563blackghost):

you are going to want to change this into a slope-intercept form so y = mx + b if you dont know how this link might help http://www.mathwarehouse.com/algebra/linear_equation/standard-form-to-slope-intercept-form.php

OpenStudy (texaschic101):

4x + y = -3 y = -4x - 3 In y = mx + b form, the slope is in the m position. If the slope is negative, then the line is decreasing

OpenStudy (anonymous):

So falling?